Two waves travel in the same medium at the same speed. One has wavelength 0.0382 m and frequency 9.67×106 Hz. The other has wavelength 0.0481 m. What is the period of the second wave?
answer in seconds.
1
Expert's answer
2011-04-06T13:31:16-0400
If the speed of waves is the same for each of it, λ1f1 = λ2f2.
0.0382*9.67×106 = 0.0481*f2,
f2 = 7.68×106 Hz.